OpenStudy (anonymous):

Between 1 /5 and 1/ 4 of the students in a class received an A on a history test. Which decimal is between 1/ 5 and 1/ 4 ? A 0.17 B 0.23 C 0.26 D 0.32

jimthompson5910 (jim_thompson5910):

Use a calculator to find that 1/5 = 0.20 1/4 = 0.25 So you need to find a number between 0.20 and 0.25

OpenStudy (anonymous):

1/5=.20 1/7=.25 so in beween them is .23 so the answer is B